# TokBench: Evaluating Your Visual Tokenizer before Visual Generation

### v1

@@ Abstract
In this work, we reveal the limitations of visual tokenizers and VAEs in preserving fine-grained features, and propose a benchmark to evaluate reconstruction performance for two challenging visual contents: text and face. Image tokenization has significantly advanced visual generation and multimodal modeling, particularly with autoregressive models due to the modeling simplicity of discrete tokens. Autoregressive models typically rely on image tokenizers to compress images into discrete tokens for sequential prediction, whereas diffusion models often operate on continuous latent space to reduce computational costs. However, both visual compression approaches inevitably lose visual information, thereby limiting the upper bound of visual generation quality. To evaluate how these compression losses affect text and faces, the most human-sensitive visual elements, we first collect and curate a collection of text and faces images from existing datasets, ensuring clarity and diversity. For text reconstruction, we employ OCR models to assess the recognition accuracy of the reconstructed text, and then we measure feature similarity between original and reconstructed faces thereby quantifying faces reconstruction fidelity. Our method is highly lightweight, requiring just 2GB memory and 4 minutes to complete evaluations. With our benchmark, we analyze the reconstruction quality of text and faces at various scales across different image tokenizers and VAEs. Our results demonstrate that modern visual tokenizers still struggle to preserve fine-grained features, particularly at smaller scales. Furthermore, we extend this evaluation framework to the video, conducting a comprehensive analysis of video tokenizers. Additionally, we find that traditional metrics fail to accurately reflect the reconstruction performance for faces and text, while our proposed metrics serve as an effective complement.
@@ 4 Experiments
Resize 86.05 93.02 92.98 96.53 0.85 0.93 5.39 0.06 27.71 0.84
Discrete VQGAN esser2021taming 0.05 1.10 4.34 8.22 0.05 0.10 12.63 0.36 17.29 0.55
Chameleon team2024chameleon 0.11 2.87 4.67 12.08 0.08 0.18 17.32 0.36 17.81 0.56
LlamaGen llamagen 0.16 4.28 5.41 14.77 0.07 0.15 11.17 0.30 18.22 0.58
VAR var 1.24 15.74 10.89 34.19 0.10 0.23 8.91 0.24 19.98 0.63
TiTok titok 0.05 0.09 3.04 4.23 0.03 0.04 16.25 0.52 13.54 0.47
FlexTok flextok 0.55 6.95 7.80 21.09 0.06 0.15 8.87 0.35 17.37 0.57
MaskBit weber2024maskbit 0.16 2.54 4.45 10.85 0.06 0.11 12.53 0.38 18.07 0.57
OmniTokenizer OmniTokenizer 2.14 20.63 13.24 39.14 0.15 0.37 9.26 0.30 15.15 0.59
TokenFlow tokenflow 0.28 6.73 6.41 20.46 0.07 0.15 9.09 0.28 18.74 0.59
UniTok ma2025unitok 13.53 44.59 38.73 65.84 0.15 0.35 7.82 0.20 21.15 0.66
O-MAGVIT2 openmagvit2 0.34 7.52 6.46 20.99 0.08 0.19 8.51 0.27 19.05 0.60
O-MAGVIT2(pretrain) openmagvit2 0.80 10.58 9.59 27.59 0.08 0.20 8.39 0.27 19.33 0.61
LlamaGen(F8) llamagen 4.39 29.41 19.69 49.00 0.17 0.40 8.65 0.19 21.50 0.67
O-MAGVIT2(F8) openmagvit2 9.33 40.24 30.82 59.97 0.23 0.48 7.88 0.17 22.53 0.70
Continuous DC-AE xie2024sana 1.42 16.35 10.95 33.82 0.10 0.26 12.88 0.23 20.88 0.65
VA-VAE yao2025vavae 6.92 37.04 25.14 56.32 0.22 0.49 6.68 0.16 22.94 0.70
SD-XL sdxl 6.94 34.21 25.03 53.68 0.18 0.42 7.60 0.19 22.52 0.69
SD-3.5 sd3 36.26 67.04 59.04 80.58 0.43 0.70 7.11 0.13 24.89 0.75
FLUX.1-dev flux2024 50.69 75.91 70.70 86.42 0.52 0.76 6.42 0.11 25.50 0.77
Table 1: Performance of discrete and continuous tokenizer on TokBench.s’ and ‘m’denote the average metrics for small-scale instances and all instances, respectively. In this table, we compute traditional metrics such as rFID across both the text set and face set.
@@ 4.1 Evaluation Setting
Resize 86.05 94.65 98.37 93.02 92.98 97.22 99.38 96.53 5.66 0.07 25.40 0.81
Discrete VQGAN 0.05 0.12 3.14 1.10 4.34 5.33 15.00 8.22 15.66 0.33 17.17 0.58
Chameleon 0.11 0.31 8.19 2.87 4.67 6.65 24.91 12.08 17.60 0.33 17.66 0.59
LlamaGen 0.16 0.44 12.25 4.28 5.41 7.50 31.40 14.77 14.23 0.29 18.04 0.61
VAR 1.24 6.72 39.26 15.74 10.89 26.26 65.42 34.19 10.30 0.22 19.74 0.66
TiTok 0.05 0.06 0.17 0.09 3.04 4.07 5.58 4.23 18.41 0.50 13.80 0.50
FlexTok 0.55 2.24 18.06 6.95 7.80 14.26 41.21 21.09 11.01 0.31 17.58 0.61
MaskBit 0.16 0.19 7.26 2.54 4.45 5.72 22.37 10.85 17.05 0.37 17.90 0.60
OmniTokenizer 2.14 8.46 51.28 20.63 13.24 30.50 73.67 39.14 12.70 0.30 14.73 0.62
TokenFlow 0.28 1.62 18.29 6.73 6.41 12.34 42.64 20.46 11.04 0.26 18.61 0.62
UniTok 13.53 42.87 77.35 44.59 38.73 68.51 90.27 65.84 9.21 0.19 20.58 0.68
O-MAGVIT2 0.34 1.49 20.73 7.52 6.46 12.41 44.10 20.99 10.18 0.25 18.85 0.63
O-MAGVIT2(pretrain) 0.80 3.17 27.76 10.58 9.59 19.15 54.02 27.59 9.83 0.24 19.15 0.64
LlamaGen(F8) 4.39 17.86 65.97 29.41 19.69 44.56 82.76 49.00 10.51 0.18 20.85 0.68
O-MAGVIT2(F8) 9.33 34.16 77.24 40.24 30.82 59.89 89.19 59.97 8.99 0.16 21.71 0.71
Continuous DC-AE 1.42 5.16 42.45 16.35 10.95 24.06 66.45 33.82 14.61 0.22 20.42 0.67
VA-VAE 6.92 28.25 75.96 37.04 25.14 55.30 88.52 56.32 8.22 0.16 21.94 0.71
SD-XL 6.94 24.83 70.85 34.21 25.03 50.96 85.03 53.68 8.93 0.18 21.69 0.70
SD-3.5 36.26 72.18 92.68 67.04 59.04 85.64 97.06 80.58 8.40 0.13 23.46 0.75
FLUX.1-dev 50.69 82.14 94.89 75.91 70.70 90.67 97.90 86.42 7.19 0.12 23.93 0.76
@@
Resize 92.51 98.18 98.86 96.52 96.25 99.24 99.64 98.38 0.26 0.01 29.80 0.91
Discrete VQGAN 0.15 0.76 17.45 6.12 5.20 8.99 37.77 17.32 6.87 0.19 19.24 0.65
Chameleon 0.60 2.67 31.39 11.55 7.63 17.82 54.95 26.80 5.61 0.17 19.81 0.66
LlamaGen 0.67 3.93 40.43 15.01 7.76 20.17 63.39 30.44 5.28 0.15 20.21 0.68
VAR 3.71 20.59 63.62 29.31 18.01 49.56 82.44 50.00 3.78 0.12 21.27 0.73
OmniTokenizer 6.21 38.33 82.91 42.48 23.44 65.99 92.66 60.70 5.67 0.20 15.00 0.67
TokenFlow 1.06 6.27 44.88 17.40 10.00 28.39 68.07 35.49 5.14 0.15 20.46 0.68
UniTok 17.25 51.86 81.20 50.10 44.75 76.28 92.20 71.08 3.27 0.09 22.54 0.76
O-MAGVIT2 1.40 9.51 54.04 21.65 10.79 33.15 74.94 39.63 3.65 0.13 21.11 0.71
O-MAGVIT2(pretrain) 3.02 16.25 62.72 27.33 16.87 44.50 80.48 47.28 3.51 0.12 21.54 0.72
LlamaGen(F8) 12.13 56.66 88.89 52.56 34.11 77.45 95.39 68.99 2.60 0.07 23.46 0.78
O-MAGVIT2(F8) 20.66 70.36 90.66 60.56 46.60 85.41 96.00 76.00 2.34 0.07 24.39 0.80
Continuous DC-AE 5.31 30.10 79.33 38.25 20.91 57.78 89.98 56.22 2.33 0.09 23.24 0.76
VA-VAE 12.72 58.73 88.43 53.30 34.80 78.86 95.30 69.65 2.23 0.07 24.07 0.79
SD-XL 16.53 62.87 91.20 56.86 40.43 80.83 96.40 72.55 2.00 0.06 24.67 0.80
SD-3.5 56.55 91.64 97.33 81.84 75.56 96.44 98.91 90.30 1.33 0.03 26.57 0.85
FLUX.1-dev 70.29 94.62 98.02 87.64 84.67 97.65 99.26 93.86 0.73 0.03 27.25 0.86
@@
Resize 95.15 98.39 99.30 97.61 97.97 99.33 99.77 99.02 0.18 0.01 inf 0.96
Discrete VQGAN 0.76 2.69 41.53 15.00 7.90 15.03 63.47 28.80 4.03 0.11 21.67 0.74
Chameleon 3.00 8.22 59.33 23.52 14.46 29.14 77.56 40.39 2.98 0.09 22.33 0.75
LlamaGen 3.30 10.62 67.63 27.19 14.13 33.02 83.57 43.58 3.35 0.09 22.74 0.77
VAR 9.64 30.08 75.35 38.36 29.07 59.16 89.51 59.25 4.85 0.10 22.40 0.79
OmniTokenizer 14.27 54.67 91.49 53.48 36.63 76.92 96.53 70.02 4.13 0.16 15.30 0.74
TokenFlow 4.46 14.86 68.57 29.30 18.62 41.21 84.43 48.09 3.34 0.09 23.26 0.78
UniTok 26.90 47.91 74.29 49.70 54.36 72.48 87.93 71.59 4.02 0.07 24.22 0.83
O-MAGVIT2 5.76 20.74 77.71 34.74 19.42 47.04 89.63 52.03 2.47 0.07 23.86 0.80
O-MAGVIT2(pretrain) 9.08 29.35 79.77 39.40 28.65 57.43 90.78 58.95 2.32 0.07 24.46 0.81
LlamaGen(F8) 25.42 71.63 94.61 63.89 50.33 86.45 97.95 78.24 1.74 0.04 26.57 0.86
O-MAGVIT2(F8) 35.29 78.91 94.84 69.68 60.97 90.36 98.03 83.12 2.21 0.06 27.07 0.88
Continuous DC-AE 15.32 48.36 92.72 52.14 35.47 71.69 96.89 68.02 1.11 0.04 27.01 0.85
VA-VAE 25.14 69.54 93.84 62.84 48.94 85.17 97.52 77.21 1.59 0.04 27.31 0.87
SD-XL 31.41 75.83 96.29 67.84 56.60 88.34 98.52 81.15 1.01 0.03 28.60 0.88
SD-3.5 74.88 95.76 98.50 89.71 87.57 98.15 99.44 95.05 0.54 0.02 29.80 0.92
FLUX.1-dev 83.71 96.83 98.72 93.09 92.68 98.69 99.52 96.96 0.41 0.01 30.55 0.94
@@ A.1 Tokenizer Selection
Resize 0.85 0.97 0.98 0.93 7.83 0.05 29.83 0.87
Discrete VQGAN 0.05 0.08 0.17 0.10 18.08 0.38 17.39 0.52
Chameleon 0.08 0.15 0.30 0.18 25.87 0.39 17.94 0.53
LlamaGen 0.07 0.11 0.26 0.15 15.30 0.32 18.38 0.55
VAR 0.10 0.20 0.41 0.23 13.11 0.25 20.20 0.61
TiTok 0.03 0.03 0.05 0.04 23.11 0.53 13.31 0.43
FlexTok 0.06 0.12 0.25 0.15 13.54 0.38 17.18 0.54
MaskBit 0.06 0.09 0.19 0.11 15.92 0.39 18.23 0.55
OmniTokenizer 0.15 0.34 0.61 0.37 12.06 0.31 15.53 0.56
TokenFlow 0.07 0.13 0.26 0.15 13.43 0.30 18.85 0.56
UniTok 0.15 0.32 0.58 0.35 11.25 0.21 21.66 0.65
O-MAGVIT2 0.08 0.15 0.34 0.19 12.91 0.29 19.24 0.58
O-MAGVIT2(pretrain) 0.08 0.16 0.35 0.20 12.92 0.29 19.49 0.59
LlamaGen(F8) 0.17 0.38 0.66 0.40 12.01 0.20 22.09 0.66
O-MAGVIT2(F8) 0.23 0.48 0.74 0.48 11.47 0.18 23.27 0.69
Continuous DC-AE 0.10 0.21 0.45 0.26 17.58 0.25 21.30 0.62
VA-VAE 0.22 0.48 0.76 0.49 9.26 0.16 23.85 0.70
SD-XL 0.18 0.40 0.69 0.42 11.19 0.20 23.29 0.68
SD-3.5 0.43 0.76 0.92 0.70 9.91 0.13 26.20 0.75
FLUX.1-dev 0.52 0.83 0.95 0.76 9.32 0.11 26.94 0.78
@@
Resize 0.95 0.99 1.00 0.98 0.08 0.00 37.34 0.97
Discrete VQGAN 0.08 0.11 0.37 0.19 7.33 0.23 20.42 0.61
Chameleon 0.13 0.21 0.50 0.28 6.62 0.22 20.98 0.61
LlamaGen 0.11 0.17 0.48 0.25 5.28 0.18 21.41 0.65
VAR 0.14 0.24 0.57 0.32 4.59 0.15 22.16 0.69
OmniTokenizer 0.24 0.45 0.80 0.50 5.11 0.20 15.93 0.63
TokenFlow 0.11 0.16 0.45 0.24 6.48 0.19 21.39 0.64
UniTok 0.22 0.36 0.74 0.44 3.95 0.11 24.34 0.74
O-MAGVIT2 0.13 0.22 0.58 0.31 4.67 0.16 22.40 0.67
O-MAGVIT2(pretrain) 0.13 0.22 0.57 0.31 4.55 0.16 22.66 0.68
LlamaGen(F8) 0.28 0.49 0.83 0.53 2.73 0.08 25.49 0.77
O-MAGVIT2(F8) 0.35 0.58 0.88 0.61 2.80 0.07 26.81 0.80
Continuous DC-AE 0.16 0.29 0.71 0.39 2.81 0.11 25.08 0.73
VA-VAE 0.31 0.54 0.87 0.57 2.41 0.07 26.84 0.79
SD-XL 0.29 0.51 0.87 0.55 2.46 0.07 27.14 0.79
SD-3.5 0.61 0.84 0.98 0.81 1.20 0.03 30.06 0.87
FLUX.1-dev 0.71 0.89 0.98 0.86 0.71 0.02 31.06 0.90

### v2

@@ Abstract
In this work, we reveal the limitations of visual tokenizers and VAEs in preserving fine-grained features, and propose a benchmark to evaluate reconstruction performance for two challenging visual contents: text and face. Visual tokenizers and VAEs have significantly advanced visual generation and multimodal modeling by providing more efficient compressed or quantized image representations. However, while helping production models reduce computational burdens, the information loss from image compression fundamentally limits the upper bound of visual generation quality. To evaluate this upper bound, we focus on assessing reconstructed text and facial features since they typically: 1) exist at smaller scales, 2) contain dense and rich textures, 3) are prone to collapse, and 4) are highly sensitive to human vision. We first collect and curate a diverse set of clear text and face images from existing datasets. Unlike approaches using VLM models, we employ established OCR and face recognition models for evaluation, ensuring accuracy while maintaining an exceptionally lightweight assessment process requiring just 2GB memory and 4 minutes to complete. Using our benchmark, we analyze text and face reconstruction quality across various scales for different image tokenizers and VAEs. Our results show modern visual tokenizers still struggle to preserve fine-grained features, especially at smaller scales. We further extend this evaluation framework to video, conducting comprehensive analysis of video tokenizers. Additionally, we demonstrate that traditional metrics fail to accurately reflect reconstruction performance for faces and text, while our proposed metrics serve as an effective complement.
@@ 4 Experiments
Resize 1×1$\times$1 × 86.05 93.02 92.98 96.53 0.85 0.93 5.39 0.06 27.71 0.84
Discrete TiTok titok $1D$ 0.05 0.09 3.04 4.23 0.03 0.04 16.25 0.52 13.54 0.47
FlexTok flextok $1D$ 0.55 6.95 7.80 21.09 0.06 0.15 8.87 0.35 17.37 0.57
VQGAN esser2021taming 16×16$\times$16 × 0.05 1.10 4.34 8.22 0.05 0.10 12.63 0.36 17.29 0.55
Chameleon team2024chameleon 16×16$\times$16 × 0.11 2.87 4.67 12.08 0.08 0.18 17.32 0.36 17.81 0.56
LlamaGen llamagen 16×16$\times$16 × 0.16 4.28 5.41 14.77 0.07 0.15 11.17 0.30 18.22 0.58
VAR var 16×16$\times$16 × 1.24 15.74 10.89 34.19 0.10 0.23 8.91 0.24 19.98 0.63
MaskBit weber2024maskbit 16×16$\times$16 × 0.16 2.54 4.45 10.85 0.06 0.11 12.53 0.38 18.07 0.57
TokenFlow tokenflow 16×16$\times$16 × 0.28 6.73 6.41 20.46 0.07 0.15 9.09 0.28 18.74 0.59
O-MAGVIT2 openmagvit2 16×16$\times$16 × 0.34 7.52 6.46 20.99 0.08 0.19 8.51 0.27 19.05 0.60
O-MAGVIT2(pretrain) openmagvit2 16×16$\times$16 × 0.80 10.58 9.59 27.59 0.08 0.20 8.39 0.27 19.33 0.61
UniTok ma2025unitok 16×16$\times$16 × 13.53 44.59 38.73 65.84 0.15 0.35 7.82 0.20 21.15 0.66
OmniTokenizer OmniTokenizer 8×8$\times$8 × 2.14 20.63 13.24 39.14 0.15 0.37 9.26 0.30 15.15 0.59
LlamaGen(F8) llamagen 8×8$\times$8 × 4.39 29.41 19.69 49.00 0.17 0.40 8.65 0.19 21.50 0.67
O-MAGVIT2(F8) openmagvit2 8×8$\times$8 × 9.33 40.24 30.82 59.97 0.23 0.48 7.88 0.17 22.53 0.70
Continuous DC-AE xie2024sana 32×32$\times$32 × 1.42 16.35 10.95 33.82 0.10 0.26 12.88 0.23 20.88 0.65
VA-VAE yao2025vavae 16×16$\times$16 × 6.92 37.04 25.14 56.32 0.22 0.49 6.68 0.16 22.94 0.70
SD-XL sdxl 8×8$\times$8 × 6.94 34.21 25.03 53.68 0.18 0.42 7.60 0.19 22.52 0.69
SD-3.5 sd3 8×8$\times$8 × 36.26 67.04 59.04 80.58 0.43 0.70 7.11 0.13 24.89 0.75
FLUX.1-dev flux2024 8×8$\times$8 × 50.69 75.91 70.70 86.42 0.52 0.76 6.42 0.11 25.50 0.77
Table 1: Performance of discrete and continuous tokenizer on TokBench.s’ and ‘m’denote the average metrics for small-scale instances and all scales, respectively. In this table, we compute traditional metrics such as rFID across both the text set and face set. The ‘Factor’ denotes the downsampling ratio in latent space, while ‘1D’ indicates that images are encoded into one-dimension.
@@ 4.1 Evaluation Setting
Resize 1×1$\times$1 × 86.05 94.65 98.37 93.02 92.98 97.22 99.38 96.53 5.66 0.07 25.40 0.81
Discrete TiTok $1D$ 0.05 0.06 0.17 0.09 3.04 4.07 5.58 4.23 18.41 0.50 13.80 0.50
FlexTok $1D$ 0.55 2.24 18.06 6.95 7.80 14.26 41.21 21.09 11.01 0.31 17.58 0.61
VQGAN 16×16$\times$16 × 0.05 0.12 3.14 1.10 4.34 5.33 15.00 8.22 15.66 0.33 17.17 0.58
Chameleon 16×16$\times$16 × 0.11 0.31 8.19 2.87 4.67 6.65 24.91 12.08 17.60 0.33 17.66 0.59
LlamaGen 16×16$\times$16 × 0.16 0.44 12.25 4.28 5.41 7.50 31.40 14.77 14.23 0.29 18.04 0.61
VAR 16×16$\times$16 × 1.24 6.72 39.26 15.74 10.89 26.26 65.42 34.19 10.30 0.22 19.74 0.66
MaskBit 16×16$\times$16 × 0.16 0.19 7.26 2.54 4.45 5.72 22.37 10.85 17.05 0.37 17.90 0.60
TokenFlow 16×16$\times$16 × 0.28 1.62 18.29 6.73 6.41 12.34 42.64 20.46 11.04 0.26 18.61 0.62
O-MAGVIT2 16×16$\times$16 × 0.34 1.49 20.73 7.52 6.46 12.41 44.10 20.99 10.18 0.25 18.85 0.63
O-MAGVIT2(pretrain) 16×16$\times$16 × 0.80 3.17 27.76 10.58 9.59 19.15 54.02 27.59 9.83 0.24 19.15 0.64
UniTok 16×16$\times$16 × 13.53 42.87 77.35 44.59 38.73 68.51 90.27 65.84 9.21 0.19 20.58 0.68
OmniTokenizer 8×8$\times$8 × 2.14 8.46 51.28 20.63 13.24 30.50 73.67 39.14 12.70 0.30 14.73 0.62
LlamaGen 8×8$\times$8 × 4.39 17.86 65.97 29.41 19.69 44.56 82.76 49.00 10.51 0.18 20.85 0.68
O-MAGVIT2 8×8$\times$8 × 9.33 34.16 77.24 40.24 30.82 59.89 89.19 59.97 8.99 0.16 21.71 0.71
Continuous DC-AE 32×32$\times$32 × 1.42 5.16 42.45 16.35 10.95 24.06 66.45 33.82 14.61 0.22 20.42 0.67
VA-VAE 16×16$\times$16 × 6.92 28.25 75.96 37.04 25.14 55.30 88.52 56.32 8.22 0.16 21.94 0.71
SD-XL 8×8$\times$8 × 6.94 24.83 70.85 34.21 25.03 50.96 85.03 53.68 8.93 0.18 21.69 0.70
SD-3.5 8×8$\times$8 × 36.26 72.18 92.68 67.04 59.04 85.64 97.06 80.58 8.40 0.13 23.46 0.75
FLUX.1-dev 8×8$\times$8 × 50.69 82.14 94.89 75.91 70.70 90.67 97.90 86.42 7.19 0.12 23.93 0.76
@@
Resize 1×1$\times$1 × 92.51 98.18 98.86 96.52 96.25 99.24 99.64 98.38 0.26 0.01 29.80 0.91
Discrete VQGAN 16×16$\times$16 × 0.15 0.76 17.45 6.12 5.20 8.99 37.77 17.32 6.87 0.19 19.24 0.65
Chameleon 16×16$\times$16 × 0.60 2.67 31.39 11.55 7.63 17.82 54.95 26.80 5.61 0.17 19.81 0.66
LlamaGen 16×16$\times$16 × 0.67 3.93 40.43 15.01 7.76 20.17 63.39 30.44 5.28 0.15 20.21 0.68
VAR 16×16$\times$16 × 3.71 20.59 63.62 29.31 18.01 49.56 82.44 50.00 3.78 0.12 21.27 0.73
TokenFlow 16×16$\times$16 × 1.06 6.27 44.88 17.40 10.00 28.39 68.07 35.49 5.14 0.15 20.46 0.68
O-MAGVIT2 16×16$\times$16 × 1.40 9.51 54.04 21.65 10.79 33.15 74.94 39.63 3.65 0.13 21.11 0.71
O-MAGVIT2(pretrain) 16×16$\times$16 × 3.02 16.25 62.72 27.33 16.87 44.50 80.48 47.28 3.51 0.12 21.54 0.72
UniTok 16×16$\times$16 × 17.25 51.86 81.20 50.10 44.75 76.28 92.20 71.08 3.27 0.09 22.54 0.76
OmniTokenizer 8×8$\times$8 × 6.21 38.33 82.91 42.48 23.44 65.99 92.66 60.70 5.67 0.20 15.00 0.67
LlamaGen 8×8$\times$8 × 12.13 56.66 88.89 52.56 34.11 77.45 95.39 68.99 2.60 0.07 23.46 0.78
O-MAGVIT2 8×8$\times$8 × 20.66 70.36 90.66 60.56 46.60 85.41 96.00 76.00 2.34 0.07 24.39 0.80
Continuous DC-AE 32×32$\times$32 × 5.31 30.10 79.33 38.25 20.91 57.78 89.98 56.22 2.33 0.09 23.24 0.76
VA-VAE 16×16$\times$16 × 12.72 58.73 88.43 53.30 34.80 78.86 95.30 69.65 2.23 0.07 24.07 0.79
SD-XL 8×8$\times$8 × 16.53 62.87 91.20 56.86 40.43 80.83 96.40 72.55 2.00 0.06 24.67 0.80
SD-3.5 8×8$\times$8 × 56.55 91.64 97.33 81.84 75.56 96.44 98.91 90.30 1.33 0.03 26.57 0.85
FLUX.1-dev 8×8$\times$8 × 70.29 94.62 98.02 87.64 84.67 97.65 99.26 93.86 0.73 0.03 27.25 0.86
@@
Resize 1×1$\times$1 × 95.15 98.39 99.30 97.61 97.97 99.33 99.77 99.02 0.18 0.01 inf 0.96
Discrete VQGAN 16×16$\times$16 × 0.76 2.69 41.53 15.00 7.90 15.03 63.47 28.80 4.03 0.11 21.67 0.74
Chameleon 16×16$\times$16 × 3.00 8.22 59.33 23.52 14.46 29.14 77.56 40.39 2.98 0.09 22.33 0.75
LlamaGen 16×16$\times$16 × 3.30 10.62 67.63 27.19 14.13 33.02 83.57 43.58 3.35 0.09 22.74 0.77
VAR 16×16$\times$16 × 9.64 30.08 75.35 38.36 29.07 59.16 89.51 59.25 4.85 0.10 22.40 0.79
TokenFlow 16×16$\times$16 × 4.46 14.86 68.57 29.30 18.62 41.21 84.43 48.09 3.34 0.09 23.26 0.78
O-MAGVIT2 16×16$\times$16 × 5.76 20.74 77.71 34.74 19.42 47.04 89.63 52.03 2.47 0.07 23.86 0.80
O-MAGVIT2(pretrain) 16×16$\times$16 × 9.08 29.35 79.77 39.40 28.65 57.43 90.78 58.95 2.32 0.07 24.46 0.81
UniTok 16×16$\times$16 × 26.90 47.91 74.29 49.70 54.36 72.48 87.93 71.59 4.02 0.07 24.22 0.83
OmniTokenizer 8×8$\times$8 × 14.27 54.67 91.49 53.48 36.63 76.92 96.53 70.02 4.13 0.16 15.30 0.74
LlamaGen 8×8$\times$8 × 25.42 71.63 94.61 63.89 50.33 86.45 97.95 78.24 1.74 0.04 26.57 0.86
O-MAGVIT2 8×8$\times$8 × 35.29 78.91 94.84 69.68 60.97 90.36 98.03 83.12 2.21 0.06 27.07 0.88
Continuous DC-AE 32×32$\times$32 × 15.32 48.36 92.72 52.14 35.47 71.69 96.89 68.02 1.11 0.04 27.01 0.85
VA-VAE 16×16$\times$16 × 25.14 69.54 93.84 62.84 48.94 85.17 97.52 77.21 1.59 0.04 27.31 0.87
SD-XL 8×8$\times$8 × 31.41 75.83 96.29 67.84 56.60 88.34 98.52 81.15 1.01 0.03 28.60 0.88
SD-3.5 8×8$\times$8 × 74.88 95.76 98.50 89.71 87.57 98.15 99.44 95.05 0.54 0.02 29.80 0.92
FLUX.1-dev 8×8$\times$8 × 83.71 96.83 98.72 93.09 92.68 98.69 99.52 96.96 0.41 0.01 30.55 0.94
@@ A.1 Tokenizer Selection
Resize 1×1$\times$1 × 0.85 0.97 0.98 0.93 7.83 0.05 29.83 0.87
Discrete TiTok $1D$ 0.03 0.03 0.05 0.04 23.11 0.53 13.31 0.43
FlexTok $1D$ 0.06 0.12 0.25 0.15 13.54 0.38 17.18 0.54
VQGAN 16×16$\times$16 × 0.05 0.08 0.17 0.10 18.08 0.38 17.39 0.52
Chameleon 16×16$\times$16 × 0.08 0.15 0.30 0.18 25.87 0.39 17.94 0.53
LlamaGen 16×16$\times$16 × 0.07 0.11 0.26 0.15 15.30 0.32 18.38 0.55
VAR 16×16$\times$16 × 0.10 0.20 0.41 0.23 13.11 0.25 20.20 0.61
MaskBit 16×16$\times$16 × 0.06 0.09 0.19 0.11 15.92 0.39 18.23 0.55
TokenFlow 16×16$\times$16 × 0.07 0.13 0.26 0.15 13.43 0.30 18.85 0.56
O-MAGVIT2 16×16$\times$16 × 0.08 0.15 0.34 0.19 12.91 0.29 19.24 0.58
O-MAGVIT2(pretrain) 16×16$\times$16 × 0.08 0.16 0.35 0.20 12.92 0.29 19.49 0.59
UniTok 16×16$\times$16 × 0.15 0.32 0.58 0.35 11.25 0.21 21.66 0.65
OmniTokenizer 8×8$\times$8 × 0.15 0.34 0.61 0.37 12.06 0.31 15.53 0.56
LlamaGen 8×8$\times$8 × 0.17 0.38 0.66 0.40 12.01 0.20 22.09 0.66
O-MAGVIT2 8×8$\times$8 × 0.23 0.48 0.74 0.48 11.47 0.18 23.27 0.69
Continuous DC-AE 32×32$\times$32 × 0.10 0.21 0.45 0.26 17.58 0.25 21.30 0.62
VA-VAE 16×16$\times$16 × 0.22 0.48 0.76 0.49 9.26 0.16 23.85 0.70
SD-XL 8×8$\times$8 × 0.18 0.40 0.69 0.42 11.19 0.20 23.29 0.68
SD-3.5 8×8$\times$8 × 0.43 0.76 0.92 0.70 9.91 0.13 26.20 0.75
FLUX.1-dev 8×8$\times$8 × 0.52 0.83 0.95 0.76 9.32 0.11 26.94 0.78
@@
Resize 1×1$\times$1 × 0.95 0.99 1.00 0.98 0.08 0.00 37.34 0.97
Discrete VQGAN 16×16$\times$16 × 0.08 0.11 0.37 0.19 7.33 0.23 20.42 0.61
Chameleon 16×16$\times$16 × 0.13 0.21 0.50 0.28 6.62 0.22 20.98 0.61
LlamaGen 16×16$\times$16 × 0.11 0.17 0.48 0.25 5.28 0.18 21.41 0.65
VAR 16×16$\times$16 × 0.14 0.24 0.57 0.32 4.59 0.15 22.16 0.69
TokenFlow 16×16$\times$16 × 0.11 0.16 0.45 0.24 6.48 0.19 21.39 0.64
O-MAGVIT2 16×16$\times$16 × 0.13 0.22 0.58 0.31 4.67 0.16 22.40 0.67
O-MAGVIT2(pretrain) 16×16$\times$16 × 0.13 0.22 0.57 0.31 4.55 0.16 22.66 0.68
UniTok 16×16$\times$16 × 0.22 0.36 0.74 0.44 3.95 0.11 24.34 0.74
OmniTokenizer 8×8$\times$8 × 0.24 0.45 0.80 0.50 5.11 0.20 15.93 0.63
LlamaGen 8×8$\times$8 × 0.28 0.49 0.83 0.53 2.73 0.08 25.49 0.77
O-MAGVIT2 8×8$\times$8 × 0.35 0.58 0.88 0.61 2.80 0.07 26.81 0.80
Continuous DC-AE 32×32$\times$32 × 0.16 0.29 0.71 0.39 2.81 0.11 25.08 0.73
VA-VAE 16×16$\times$16 × 0.31 0.54 0.87 0.57 2.41 0.07 26.84 0.79
SD-XL 8×8$\times$8 × 0.29 0.51 0.87 0.55 2.46 0.07 27.14 0.79
SD-3.5 8×8$\times$8 × 0.61 0.84 0.98 0.81 1.20 0.03 30.06 0.87
FLUX.1-dev 8×8$\times$8 × 0.71 0.89 0.98 0.86 0.71 0.02 31.06 0.90